1.    Suppose the following information is true for country A:

 

a.    Average annual income :                                 $6,500 per family

b.    Average consumption of meat:                        200 pounds

c.    Estimated income elasticity of demand:          10

d.    Average price per pound of meat:          $1.00

e.    Population composed of one million families

Let us assume that the government wants to provide adequate nutrition to the families so that each family is able to afford 250 pounds of meat per year. This can be done by transferring income to the families by the government.

By using the above information, compute the cost of this program.


1
Expert's answer
2022-03-01T14:50:52-0500

Computation of Cost of Program:

Given,

Original consumption = 200 pounds

New consumption = 250 pounds

Change in consumption= 250-200 = 50 pounds

So,

"\% \Delta\ in\ consumption" = "\frac{50}{200} \times100 = 25\%"

 

Given,

Original Income = 6500

Income elasticity of demand = 10

"Income\ Elasticity\ of\ demand = \frac{\%\Delta\ in\ consumption}{\%\Delta\ in\ Income}"

"10 = \frac{25\%} {\%\ change\ in\ income}"

"\% change in income = 2.5\%"

 

Increase in Income required = "6500\times2.50\%" = $162.50 per family

Number of families = 1000,000 families

So,

Cost of Program = "(1000,000) \times(162.50 )= \$ 162,500,000"